OpenStudy (anonymous):

sum of sequence using gauss approach 1+2+3+4+..101

OpenStudy (anonymous):

I don't know what gauss approach is, just give you idea note that 1 +2 +3 +4 + 5+..................+95 + 96 + 97 + 98 + 99 , let 100 and 101 aside now, pick numbers from leftside and rightside in order : 1+99 =100 2 + 98 =100 3 + 97 =100 4 + 96 = 100 5 + 95 = 100 so on and so on. you have 50 pair of that kind of number to count. that means you have 50*100 = 5000 now add 100 and 101 in, I have 5201 is the final answer.